# Generalization of Gibbs and Langevin Monte Carlo Algorithms in the Interpolation Regime

### v1

@@ B.1 Sub-gaussian losses
By the bound on $\ell$ we have
$\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\leq\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}\text{.}$ (8)
(i) From Markov’s inequality we have (in analogy to the proof of Theorem 3.2) with probability at least $1-\delta$ as $x\sim\mu^{n}$ and $h\sim\nu_{\beta}\left(\mathbf{x}\right)$, that
$\leq$ $\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]+\ln\left(1/\delta\right)$
$\leq$ $\ln\left(\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]+e^{M+\beta m}\epsilon_{\beta}\right)+\ln\left(1/\delta\right)$
$=$ $\ln\left(\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+e^{M+\beta m}\epsilon_{\beta}\right)+\ln\left(1/\delta\right)$
$\leq$ $\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\left(2e^{M+\beta m}\epsilon_{\beta}\right)+\ln\left(1/\delta\right).$
In the second inequality we used $\ln Z_{\beta}\left(\mathbf{x}\right)\leq 0$ and in the last line we used for $a,b\geq 1$ that $\ln\left(a+b\right)\leq\ln\max\left\{a,b\right\}+\ln 2\leq\ln a+\ln b+\ln 2=\ln a+\ln 2b$. Subtract $\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$, use (4) and (8).
(ii) Again with Markov’s inequality, with probability at least $1-\delta$ as $x\sim\mu^{n}$
$\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)\right]$
$\leq$ $\ln\mathbb{E}_{\mathbf{x}\sim\mu^{n}}\left[e^{\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)\right]}\right]$
$\leq$ $\ln\mathbb{E}_{\mathbf{x}\sim\mu^{n}}\left[e^{\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)\right]+\left(M+\beta m\right)\epsilon_{\beta}}\right].$
In the second inequality we used the fact that $\ln Z_{\beta}\left(\mathbf{x}\right)\leq 0$ and the bounds on $F$ and $\ell$. Then, Jensen’s inequality bounds the last line as
$\ln\mathbb{E}_{\mathbf{x}\sim\mu^{n}}\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]e^{\left(M+\beta m\right)\epsilon_{\beta}}$
$=$ $\ln\mathbb{E}_{\mathbf{x}\sim\mu^{n}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\left(M+\beta m\right)\epsilon_{\beta}.$
Again subtract $\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$, use (4) and (8). ∎
B.3 Miscellaneous lemmata
Lemma B.5.
For $0<\beta<\infty$
$\max\left\{KL\left(G_{\beta},G_{2\beta}\right),KL\left(G_{2\beta},G_{2}\right)\right\}\leq\beta\left(\mathbb{E}_{h\sim G_{\beta}}\left[\hat{L}\left(h\right)\right]-\mathbb{E}_{h\sim G_{2\beta}}\left[\hat{L}\left(h\right)\right]\right)$
Proof.
Using Lemma 3.1
@@ B.3 Miscellaneous lemmata
$\leq$ $\beta\left(\mathbb{E}_{h\sim G_{\beta}}\left[\hat{L}\left(h\right)\right]-\mathbb{E}_{h\sim G_{2\beta}}\left[\hat{L}\left(h\right)\right]\right).$
Similarly
$KL\left(G_{2\beta},G_{2}\right)$ $=$ $\mathbb{E}_{h\sim G_{2\beta}}\left[-2\beta\hat{L}\left(h\right)-\ln Z_{2\beta}+\beta\hat{L}\left(h\right)+\ln Z_{\beta}\right]$
@@
$\leq$ $\beta\left(\mathbb{E}_{h\sim G_{\beta}}\left[\hat{L}\left(h\right)\right]-\mathbb{E}_{h\sim G_{2\beta}}\left[\hat{L}\left(h\right)\right]\right).$
Appendix C Experimental details and additional results
C.1 Experimental details
All the codes to reproduce the results are provided through this https://anonymous.4open.science/r/Gibbs-Generalization-45F1. For all the experiments we use the isotropic Gaussian prior with $\mu=0$ and $\sigma=5$. This induces an L2-regularization term in the energy function that is stated in the proof of Corollary 4.2. The confidence parameter $\delta$ appearing in our bounds is set to $0.01$ for all experiments
@@ C.1 Experimental details
C.1.1 Network architecture
The fully connected networks consist of one, two, three hidden layers, each containing a constant number of units. Besides that we are using LeNet-5 architecture for MNIST and VGG16 architecture for CIFAR-10 to achieve low test error. For loss function $\ell$, we are using either bounded binary cross-entropy (BBCE) as described in Appendix D of Dziugaite & Roy (2018) or the savage loss (Masnadi-Shirazi & Vasconcelos, 2008).
@@ C.1.1 Network architecture
VGG-16 is a widely used deep convolutional neural network architecture known for its simplicity and strong performance in image classification tasks (Simonyan & Zisserman, 2014). The architecture follows a consistent design using only $3\times 3$ convolutional filters and $2\times 2$ max pooling operations throughout the network. In our implementation, VGG-16 is adapted to handle CIFAR-10’s smaller $32\times 32$ RGB images. The network consists of 13 convolutional layers organized into five blocks: the first two blocks contain two convolutional layers each with 64 and 128 filters, respectively, while the last three blocks contain three convolutional layers each with 256, 512, and 512 filters, respectively. Each block is followed by a $2\times 2$ max pooling layer for spatial downsampling. All convolutional layers employ $3\times 3$ kernels with padding to preserve spatial dimensions, and ReLU activation functions introduce non-linearity.The convolutional feature extractor is followed by a classifier head consisting of three fully connected layers: two hidden layers with 1024 neurons each, using ReLU activation, and a final output layer with 1 neurons for binary classification. We also removed dropout to ensure that SGLD minimizes the defined energy function without any additional terms.
@@
C.1.2 Minibatches
When using SGLD, we adopt minibatches of size proportional to $\sqrt{n}$. Thus, for $n=2000$ the mini-batch size is 50, and for $n=8000$ it is 100.
C.1.3 Moving average filters
@@
with $\mathbb{M}_{0}=0$ and small $\alpha$. We use the first (symmetric) form in the experiments with ULA, and the second (standard exponential moving average) form with SGLD for convenience. We set different values of $\alpha$ for the two roles: $\alpha=0.0025$ for the stopping criterion ($\mathbb{M}_{\mathrm{stop}}$) and $\alpha=0.01$ for approximating the ergodic mean ($\mathbb{M}_{\mathrm{erg}}$). The stopping rule is triggered when
@@
C.2 Experimental results
C.2.1 Illustration of Bound Computation
In this section, we demonstrate again the figure in the main body in more details. The figure 2 illustrates how our bounds are computed. The sequence of mean training losses in $\ell$ is used to compute for each $\beta$ the functional $\Gamma$ and the "KL-Bound", which corresponds to the right hand side of the inequalities in Corollary 3.3. Our bound on the test loss is then computed by applying the function $\kappa^{-1}$ to the empirical 0-1 error and to this kl-bound. The graph of "KL(Train, Test)" corresponds to the left hand side in Corollary 3.3.
It is remarkable that the close fit of the upper bound on the random labels is achieved by the adjustment of a single calibration parameter.
@@ C.2.2 Single-draws
Figure 3: SGLD on MNIST and CIFAR-10 with 2000 training examples using BBCE loss function. The first row corresponds to MNIST and the second row to CIFAR-10. Random labels are shown on the left, correct labels on the right. Both random and true labels are trained with exactly the same algorithm and parameters on a fully connected ReLU network with two hidden layers of 1000 (respectively 1500) units. The calibration factor for MNIST is 0.19, for CIFAR-10 0.22. Train error, test error and our bound for a single-draw of the 0-1 loss are plotted against $\beta$.
C.2.3 Different Architectures
In this section, we evaluate the performance of different models and architectures on both MNIST and CIFAR-10, demonstrating that our bound can be used to guide model selection. In addition to the two-hidden-layer neural networks described in Section 1, we consider fully connected neural networks with three hidden layers, containing 500 and 1000 units for MNIST and CIFAR-10, respectively. Furthermore, we employ the LeNet-5 architecture for MNIST and VGG-16 for CIFAR-10 to achieve high test accuracy. Detailed descriptions of these architectures are provided in Section C.1.1.
@@ C.2.4 ULA
C.2.5 Savage Loss function
@@
C.2.6 Uncalibrated Bounds
We show uncalibrated bounds for the MNIST dataset with the BBCE and Savage loss functions under the experimental conditions of Sections C.2.4 and C.2.5. The bounds are somewhat looser than the calibrated ones, but still far from trivial. As in all other scenarios the test errors are upper bounded correctly.
Figure 7: ULA on MNIST with 2000 training examples using BBCE and savage loss functions. The first row corresponds to MNIST with BBCE and the second row to savage. Random labels are shown on the left, correct labels on the right. Both random and true labels are trained with the same algorithm and parameters on a fully connected ReLU network with one hidden layers of 500 units. Train error, test error and our bound for the Gibbs posterior average of the 0-1 loss are plotted against $\beta$ without any calibration.
C.2.7 Real-world use cases
@@
Test Error with SGD 0.0364 0.0363 0.0308
Test Error at $\beta=64k$ 0.0498 0.0549 0.0317
Test Bound at $\beta=64k$ 0.0860 0.1314 0.0375
@@
Test Error with SGD 0.1423 0.1415 0.0933
Test Error at $\beta=64k$ 0.1719 0.1782 0.0903
Test Bound at $\beta=64k$ 0.2266 0.2807 0.2030

### v2

@@ C.5 Miscellaneous Lemmata
$\leq$ $\beta\left(\mathbb{E}_{h\sim G_{\beta}}\left[\hat{L}\left(h\right)\right]-\mathbb{E}_{h\sim G_{2\beta}}\left[\hat{L}\left(h\right)\right]\right)$
Similarly,
$KL\left(G_{2\beta},G_{\beta}\right)$ $=$ $\mathbb{E}_{h\sim G_{2\beta}}\left[-2\beta\hat{L}\left(h\right)-\ln Z_{2\beta}+\beta\hat{L}\left(h\right)+\ln Z_{\beta}\right]$
@@
$\leq$ $\beta\left(\mathbb{E}_{h\sim G_{\beta}}\left[\hat{L}\left(h\right)\right]-\mathbb{E}_{h\sim G_{2\beta}}\left[\hat{L}\left(h\right)\right]\right)$
C.6 The Calibration Factor
Let $\bar{A}=-\beta\mathbb{E}_{h\sim\nu_{\beta}}\left[\hat{L}\left(h,\mathbf{\bar{x}}\right)\right]+\Gamma\left(\nu_{0}^{K-1},\mathbf{\bar{x}},\beta_{0}^{K}\right)$ be the area estimate obtained for the random labels (corrected by $\beta\mathbb{E}_{h\sim\nu_{\beta}}\left[\hat{L}\left(h,\mathbf{\bar{x}}\right)\right]$, which should play little role for large $\beta>n$). Then
$r\left(\mathbf{x}\right)=\min\left\{r:\forall k\in\left[K\right],\text{ }\kappa^{-1}\left(\mathbb{E}_{h\sim\nu_{\beta_{k}}\left(\mathbf{x}\right)}\left[\hat{L}_{01}\left(h,\mathbf{\bar{x}}\right)\right],\frac{1}{n}\left(r\bar{A}+\ln\frac{2\sqrt{n}}{\delta}\right)\right)\geq\frac{1}{2}\right\},$
where $\mathbf{\bar{x}}$ is the training set $\mathbf{x}$ with random labels and $\hat{L}_{01}$ the empirical 01-error. The calibration value $r$ is thus the smallest factor of $\bar{A}$, for which we obtain a correct upper bound on the 01-error with random labels for all the $\beta_{k}$.
We emphasize that the calibration procedure depends only on the training data.
Appendix D Experimental Details and Additional Results
D.1 Experimental Details
All the codes to reproduce the results are provided through this https://anonymous.4open.science/r/Gibbs-Generalization-45F1. For all the experiments we use an isotropic Gaussian prior with $\mu=0$, for bounded loss with $\sigma=5$ and for unbounded loss with $\sigma=0.1$. This induces an L2-regularization term in the energy function that is stated in the proof of Corollary C.2. The confidence parameter $\delta$ appearing in our bounds is set to $0.01$ for all experiments
@@ D.1 Experimental Details
D.1.1 Networks Architecture
The fully connected networks consist of one, two, three hidden layers, each containing a constant number of units. Besides that we are using LeNet-5 architecture for MNIST and VGG16 architecture for CIFAR-10 to achieve low test error. For loss function $\ell$, we are mostly using bounded loss function such as bounded binary cross-entropy (BBCE) as described in Appendix D of Dziugaite and Roy (2018) or the Savage loss (Masnadi-Shirazi and Vasconcelos, 2008). As unbounded loss function we tried binary cross-entropy (BCE) (Section D.2.6), but with a smaller value of $\sigma$, so as to avoid excessive training errors for small values of $\beta$.
@@ D.1.1 Networks Architecture
VGG-16 is a widely used deep convolutional neural network architecture known for its simplicity and strong performance in image classification tasks (Simonyan and Zisserman, 2014). The architecture follows a consistent design using only $3\times 3$ convolutional filters and $2\times 2$ max pooling operations throughout the network. In our implementation, VGG-16 is adapted to handle CIFAR-10’s smaller $32\times 32$ RGB images. The network consists of 13 convolutional layers organized into five blocks: the first two blocks contain two convolutional layers each with 64 and 128 filters, respectively, while the last three blocks contain three convolutional layers each with 256, 512, and 512 filters, respectively. Each block is followed by a $2\times 2$ max pooling layer for spatial downsampling. All convolutional layers employ $3\times 3$ kernels with padding to preserve spatial dimensions, and ReLU activation functions introduce non-linearity.The convolutional feature extractor is followed by a classifier head consisting of three fully connected layers: two hidden layers with 1024 neurons each, using ReLU activation, and a final output layer with 1 neuron for binary classification. We also removed dropout to ensure that SGLD minimizes the defined energy function without any additional terms.
@@
D.1.2 Minibatches
When using SGLD, we adopt minibatches of size proportional to $\sqrt{n}$. Thus, for $n=2000$ the minibatch size is 50, and for $n=8000$ it is 100.
D.1.3 Moving Average Filters
@@
with $\mathbb{M}_{0}=1$ and small $\alpha$. We use the first (symmetric) form in the experiments with ULA, and the second (standard exponential moving average) form with SGLD for convenience. We set different values of $\alpha$ for the two roles: $\alpha=0.0025$ for the stopping criterion ($\mathbb{M}_{\mathrm{stop}}$) and $\alpha=0.01$ for approximating the ergodic mean ($\mathbb{M}_{\mathrm{erg}}$). The stopping rule is triggered when
@@
D.2 Experimental Results
D.2.1 Illustration of Bound Computation
In this section, we discuss figure 4 in the main body in more details. Figure 5 illustrates how our bounds are computed. The sequence of mean training losses in $\ell$ is used to compute for each $\beta$ the functional $\Gamma$ and the "KL-Bound", which corresponds to the right-hand side of the inequalities in Corollary 11. Our bound on the test loss is then computed by applying the function $\kappa^{-1}$ to the empirical 0-1 error and to this kl-bound. The graph of "KL(Train, Test)" corresponds to the left-hand side in Corollary 11. It is remarkable that the close fit of the upper bound on the random labels is achieved by the adjustment of a single calibration parameter.
@@ D.2.2 Single-draws
Figure 6: SGLD on MNIST and CIFAR-10 with 8000 training examples using BBCE loss function. The first row corresponds to MNIST and the second row to CIFAR-10. Random labels are shown on the left, correct labels on the right. Both random and true labels are trained with exactly the same algorithm and parameters on a fully connected ReLU network with two hidden layers of 1000 (respectively 1500) units. The calibration factor for MNIST is 0.77, for CIFAR-10 0.89. Train error, test error and our bound for a single-draw of the 0-1 loss are plotted against $\beta$.
D.2.3 Different Architectures
In this section, we evaluate the performance of different models and architectures on both MNIST and CIFAR-10, demonstrating that our bound can be used to guide model selection. In addition to the two-hidden-layer neural networks described in Section 4, we consider fully connected neural networks with three hidden layers, containing 500 and 1000 units for MNIST and CIFAR-10, respectively. Furthermore, we employ the LeNet-5 architecture for MNIST and VGG-16 for CIFAR-10 to achieve high test accuracy. Detailed descriptions of these architectures are provided in Section D.1.1.
@@ D.2.4 ULA
D.2.5 Savage Loss Function
@@
D.2.6 Unbounded Loss Function
In this section, we use the binary cross-entropy loss to compute the $\Gamma$ functional. Since binary cross-entropy is unbounded, the loss can become very large at high temperatures. To avoid this issue, we set the standard deviation of the Gaussian prior to 0.1 in this section. The following plot shows the results under the same setup as Section D.2.5, except that we use binary cross-entropy instead of the Savage loss.
Figure 10: ULA on MNIST with 2000 training examples using binary cross entropy loss function. Random labels are shown on the left, correct labels on the right. Both random and true labels are trained with the same algorithm and parameters on a fully connected ReLU network with one hidden layers of 500 units. The calibration factor is 0.34. Train error, test error and our bound for the Gibbs posterior average of the 0-1 loss are plotted against $\beta$.
D.2.7 Real-World Use Cases
@@
Test Error, SGD 0.0364 0.0363 0.0308
Test Error, SGLD ($\beta=64k$) 0.0498 0.0549 0.0317
Test Bound, SGLD ($\beta=64k$) 0.0860 0.1314 0.0375
@@
Test Error, SGD 0.1423 0.1415 0.0933
Test Error, SGLD ($\beta=64k$) 0.1719 0.1782 0.0903
Test Bound, SGLD ($\beta=64k$) 0.2266 0.2807 0.2030